OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 + 4x + 7

jimthompson5910 (jim_thompson5910):

Complete the square for the x terms y = x^2 + 4x + 7 y + 4 = x^2 + 4x + 7 + 4 ... see note below y + 4 = (x^2 + 4x + 4) + 7 y + 4 = (x+2)^2 + 7 y = (x+2)^2 + 7-4 y = (x+2)^2 + 3 Note: you take half of the x coordinate 4 to get 2, then you square 2 to get 4. Then you add this to both sides ----------------------------- So y = x^2 + 4x + 7 turns into y = (x+2)^2 + 3 which is now in vertex form

jimthompson5910 (jim_thompson5910):

y = (x+2)^2 + 3 is in the form y = a(x-h)^2 + k where a = 1 h = -2 k = 3 So the vertex is (-2, 3)